OpenStudy (anonymous):

the perimeter of a doubles tennis court is 76 yards. the length of the court is 2 yards more than twice its width. what is the length?

OpenStudy (anonymous):

Solve for l using substitution: 2l+2w=76 l=2w+2

OpenStudy (anonymous):

Let width be x. Thus length is 2x+2. Perimeter = 76= 2(x) + 2 ( 2x+2) 76= 6x +4 6x= 72 x= 12 Length = 2x +2 (substitute x= 12) length = 2(12) +2= 26 yards

OpenStudy (anonymous):

i don't want to. that's why i posted it on here

OpenStudy (anonymous):

thanks